OpenStudy (anonymous):

limit of 4-sqr.root(x)/ (x-16).... x is appoaching 16

OpenStudy (anonymous):

\[\lim_{x\rightarrow 16}\frac{4-\sqrt{x}}{x-16}\]

OpenStudy (anonymous):

rationalize the numerator by multiplying top and bottom by \[4+\sqrt{x}\]

OpenStudy (anonymous):

\[\frac{4-\sqrt{x}}{x-16}\times \frac{4+\sqrt{x}}{4+\sqrt{x}}\] \[=\frac{16-x}{(x-16)(4+\sqrt{x})}\] \[=\frac{-1}{x+\sqrt{16}}\]

OpenStudy (anonymous):

then replace x by 16 to get \[\frac{-1}{8}\]

OpenStudy (anonymous):

this line was a typo \[=\frac{-1}{x+\sqrt{16}}\] it should have been \[\frac{-1}{4+\sqrt{x}}\]